A grinding wheel is in the form of a uniform solid disk of radius 6.98 cm and mass 2.04 kg. It starts from rest and accelerates uniformly under the action of the constant torque of 0.599 N · m that the motor exerts on the wheel.

(a) How long does the wheel take to reach its final operating speed of 1 230 rev/min?
(b) Through how many revolutions does it turn while accelerating?
Answer #1

Moment of inertia of solid disk, I = 0.5mr2 = 0.5*2.04*(0.0698)2 = 0.005 kg-m2

Given, initial angular velocity, wo = 0 rad/s

Torque, T = 0.599 N-m

a) Given, final angular velocity, w = 1.230 rev/min = 0.128 rad/s

we know that, T = I

hence, = T/I = 0.599/0.005 = 119.8 rad/s2

using,

w = wo + *t

0.128 = 0 + 119.8*t

t = 0.00106 s

b) using,

= wot + 0.5*t2 = 0*0.00106 + 0.5*119.8*0.001062 = 6.73*10-5 radians